# Marginalia to a Theorem of Asperó and Schindler

### v2

@@ 1 Introduction
The proof systems for the finitary propositional calculus generalize in a natural way to the infinitary propositional formulas. For example, one can proceed in a way described in [§1]farah2021extender. The natural notion of coherence for an infinitary propositional formula $\phi$ then becomes “ϕ⊬not-provesitalic-$\phi\not\vdash\bot$”. It turns out that a form of completeness theorem holds in this case too, and that coherence is equivalent to consistency (cf. [Lemma 1.2]farah2021extender). By “consistency” we mean the following.
@@
In addition to the proof-theoretic characterization of consistency given by the notion of coherence, there is also a game-theoretic characterization. Recall the Model Existence Game of [Section 8.3]vaananen2011models. This game is greatly simplified when we are only interested in propositional formulas.
@@ II $w_{0}$ $w_{1}$cdots
The characterization of consistency then reads as follows.
@@
Let $\phi$ be an infinitary propositional formula in negation normal form. Then $\phi$ is consistent if and only if Player II has a winning strategy in $\Game^{\mathrm{meg}}(\phi)$.
Proof.
This follows by the proof of [Theorem 8.12]vaananen2011models. The point is that once the set of all subformulas of $\phi$ is made countable, we are in the scenario of the quoted theorem. Since the game is closed for Player II, it is generically absolute whether or not Player II wins. ∎
In this paper we are interested in the following strengthenings of the notion of consistency.
Definition 1.4.
We will describe versions of the Model Existence Game that correspond to these three notions of consistency. They are denoted by $\Game^{\mathrm{p}}_{\kappa}(o)$, $\Game^{\mathrm{sp}}_{\kappa}(o)$, and $\Game^{\mathrm{ssp}}_{\kappa}(o)$, respectively. These games depend implicitly on $\phi$ and explicitly on a beth-fixed point $\kappa$ as parameters. We will prove the following theorem.
@@
    Player II has a winning strategy in $\Game^{\mathrm{p}}_{\kappa}(o)$ (resp. $\Game^{\mathrm{sp}}_{\kappa}(o)$, $\Game^{\mathrm{ssp}}_{\kappa}(o)$), for some inaccessible $\kappa$. ∎
In addition, our methods exhibit a proper (resp. semiproper, ssp) poset that forces a model of the formula (whenever it exists). The three instances of the above theorem are discussed in detail in Sections 4, 6, and 7. In Section 2 we introduce a simple but useful notion of a problem. It allows us to recast the problem of finding a model of a formula in purely combinatorial terms. We work with this notion (instead of formulas) for the rest of the paper, except in the last section. In the penultimate section, we give an abstract sufficient condition for a problem (formula) to be ssp-consistent. This condition is essentially what is used in the proof of the main lemma of aspero2021martin. We give the exact details in the last section.
In short, we say that a formula $\phi$ is AS-good at an inaccessible $\kappa$ iff $\phi\in V_{\kappa}$, $\phi$ is consistent, and for all stationary $S\subseteq\omega_{1}$, it holds in $V^{\mathsf{Col}(\omega,<\kappa)}$ that for all $\mu\models\phi$, there exist elementary $\tau:V\longrightarrow W$ and $\hat{\mu}\models\tau(\phi)$ such that $\mathsf{crit}(\tau)=\omega_{1}^{V}\in\tau(S)$ and for all subformulas $\psi$ of $\phi$, $\hat{\mu}(\tau(\psi))=\mu(\psi)$. We will prove the following theorem.
Theorem 1.6.
Suppose $\kappa$ is inaccessible and $\phi$ is AS-good at $\kappa$. Then $\phi$ is ssp-consistent. ∎
In the last section we give an example of an AS-good formula. This example shows how to ensure the existence of embeddings that witness the AS goodness.
@@ 3 Some Simple Remarks
The side conditions in our posets will be virtual models. They were first defined by Veličković when studying alternative ways to iterate semiproper posets, see for example kasum2023iterating. In the cited paper, one can also find a detailed elaboration on how to work with virtual models. We review here only the basics. A prototype of a virtual model is an elementary submodel of some $H_{\theta}$, with $\theta$ regular, but there is more to the notion.
@@
If $M\in\mathcal{C}$ and if $\mathbb{Q}$ is a poset that belongs to $M\cap V_{\lambda_{M}}$, we can make sense of the phrase “$\mathbb{Q}$ is (semi)proper for $M$”. One way to do so is to just run the usual definition, but with respect to this more general object $M$999 The point is that $M$ is not necessarily an elementary submodel of some $H_{\theta}$. It turns out that this notion is equivalent to saying that $\mathbb{Q}$ is (semi)proper for $M\cap H_{\theta}$ for any and all regular $\theta\in(2^{2^{\kappa}},\lambda_{M})\cap M$. However, $M\cap H_{\theta}\prec H_{\theta}$, so we do not diverge much from the usual notion of (semi)properness.
@@
When passing from a virtual model $M$ to its projection $M{\downarrow}\lambda$, we lose information. In a sense, $M$ then certifies that $M{\downarrow}\lambda$ is “consistent” with more information.101010 Our intuition is that it is more complicated to embed a model into a transitive structure of a larger rank. This is not true for all models $N\in\mathcal{C}_{\lambda}$ and even if it is true for some such $N$, a virtual model $M$ such that $M{\downarrow}\lambda=N$ is not necessarily unique. We will use the following notion in order to talk about this possibility of increasing the information content of a virtual model.
@@ II $q_{-1}$ $q_{0}$
Suppose that is a proper class of inaccessible cardinals. Then the following are equivalent.
@@ 5 An Example for the Proper Consistency
The goal is to show that there exists a proper poset $\mathbb{Q}$ such that in $V^{\mathbb{Q}}$, there exists a sequence $(X_{n}:n<\omega)$ such that $X=\bigsqcup_{n<\omega}X_{n}$ and such that for all $n<\omega$, $\Gamma{\restriction}[X_{n}]^{2}=\mathrm{const}$121212 This was previously known by the results of abraham1985on. This goal can be naturally reformulated as a problem (in our formal sense).
@@
    $\phi(x)\in\mathsf{tp}^{\mathcal{H}}(a/M)$.131313 We use the notation “$\mathsf{tp}^{\mathcal{H}}(a/M)$” to denote the type of the object $a$ in the structure $\mathcal{H}$ with respect to parameters in $M$.
@@
      the first $k$ elements of $\mathsf{dom}(w_{q})$ are exactly $a_{0},\dots,a_{k-1}$,141414 We mean the first $k$ in the order $<_{\mathcal{H}}$.
@@
6 Semiproper Consistency [^15] A version of this section was done jointly with Ben De Bondt. He pursued some of these ideas in a different direction in his thesis bondt2024stationary.
@@ II $q_{-1}$ $q_{0}$
Suppose that is a proper class of inaccessible cardinals. Then the following are equivalent.
@@ 7 Stationary Set Preserving Consistency
In this section, we want to characterize the ssp consistency of $(\mathbb{H},\mathcal{D})$ in a way analogous to what we did for proper and semiproper consistency. We follow closely the outline of Sections 4 and 6. In this case, the preconditions and their ordering are exactly the same as in the semiproper case. The game $\Game^{\mathrm{ssp}}_{\kappa}(p)$ that we define below differs from the game $\Game^{\mathrm{sp}}_{\kappa}(p)$ in the questions of the second type that Player I can ask. While a semiproper poset is semiproper for club many models, this is not true for an ssp poset. What we have is that an ssp poset is semiproper for projectively stationary many models (cf. [Lemma 4.8]feng2003structure). The change in the game is such that it reflects this fact.
@@ II $q_{-1}$ $q_{0}$
Suppose that is a proper class of inaccessible cardinals. Then the following are equivalent.
@@ II $p_{-1}$ $p_{0}$ $p_{1}$cdots
The infinite plays with no rules broken are won by Player II. The set $\mathbb{C}^{\mathrm{BV}}_{\eta}$ is defined to consist of all $p\in\mathbb{P}^{\mathrm{sp}}\cap V_{\eta}$ such that Player II has a winning strategy in $\Game^{\mathrm{BV}}_{\eta}(p)$.161616 This definition is essentially due to De Bondt and Veličković.
@@ 9 An Example for the AS Goodness
In this section, we show how our methods can be used to give another proof of the main lemma of aspero2021martin. Before we precisely state what we are doing, we need a following auxiliary notion.
@@
This is shown as [Lemma 3.6]aspero2021martin. ∎
@@
  • $M{\downarrow}\lambda$, 8
  • $\Game^{\mathrm{BV}}_{\eta}(p)$, 42
@@
  • $\Game^{\mathrm{p}}_{\kappa}(p)$, 10
  • $\Game^{\mathrm{sp}}_{\kappa}(p)$, 28
  • $\Game^{\mathrm{ssp}}_{\kappa}(p)$, 36
@@
  • $\beth_{\mathrm{fix}}$, 8
  • $\delta(M)$, 7
  • $\delta_{M}$, 7
  • $\lambda_{M}$, 8
  • $\leq^{\mathrm{p}}$, 10
  • $\leq^{\mathrm{sp}}$, 28
  • $\mathbb{C}^{\mathrm{BV}}_{\eta}$, 43
  • $\mathbb{C}^{\mathrm{p}}_{\kappa}$, 11
  • $\mathbb{C}^{\mathrm{sp}}_{\kappa}$, 29
  • $\mathbb{C}^{\mathrm{ssp}}_{\kappa}$, 37
@@
  • $\mathbb{P}^{\mathrm{sp}}$, 28
  • $\mathcal{C}$, 8
  • $\mathcal{C}_{\lambda}$, 8
@@
  • $\mathsf{Hull}(M,X)$, 7
  • $\phi_{\mathrm{AS}}$, 53
  • $\theta$-well behaved, 50
  • $\vartheta(\mathcal{M})$, 51
  • $o$, 11
  • $pq$, 5
  • AS-good, 42
@@
  • problem, 5
  • proper-consistent, 3, 5
  • semiproper-consistent, 3, 5
  • ssp, 3
  • ssp-consistent, 3, 5
  • virtual model, 7
  • vm-chain, 8

### v3

@@ 1 Introduction
One of the fundamental problems in logic is to characterize when a given formula has a model and what that model is. The most important instance of this problem is the well-known Entscheidungsproblem, which asks whether a first-order theory is satisfiable. In this paper, we will be interested in certain forms of the problem which concern infinitary propositional formulas. In this case, the most basic form of a model-existence property is, of course, the consistency.
@@
A superficial inspection of this definition might cause doubts as to whether it is natural to look at models in all generic extensions. These doubts are quickly eliminated once we realize that this notion is equivalent to the property of “not proving the contradiction”. To fully appriciate this last comment, one has to recall the fact that the infinitary propositional logic comes equipped with a natural notion of a proof system, one which is a relatively straightforward generalization of the proof system for the standard (finitary) propositional logic. [^3] See, for example, [Far21, §1]. This explains the central role in Logic of this notion of consistency. The starting point of our investigation is the question of what happens with this property if we restrict the class of posets which are to be considered for adding a model to the formula.
Definition 1.2.
For each of these three notions, we will give a game-theoretic characterization. The motivation for such pursuit is the fact that the notion of consistency has, in addition to the proof-theoretic characterization, an important game theoretic one. The latter is given in terms of the Model Existence Game [^5] See [Vää11, Section 8.3]. The game presented there greatly simplifies when we are only interested in propositional formulas..
@@ II $w_{0}$ $w_{1}$cdots
The characterization of consistency that we want to imitate reads as follows.
@@
Let $\phi$ be an infinitary propositional formula in negation normal form. Then $\phi$ is consistent if and only if Player II has a winning strategy in $\Game^{\mathrm{meg}}(\phi)$. 777 This follows by the proof of [Vää11, Theorem 8.12]. The point is that once the set of all subformulas of $\phi$ is made countable, we are in the scenario of the quoted theorem. Since the game is closed for Player II, it is generically absolute whether or not Player II wins.
We will describe versions of the Model Existence Game that correspond to the notions of proper-, semi-proper, and ssp-consistency. They are denoted by $\Game^{\mathrm{p}}_{\kappa}(o)$, $\Game^{\mathrm{sp}}_{\kappa}(o)$, and $\Game^{\mathrm{ssp}}_{\kappa}(o)$ (respectively), and depend implicitly on $\phi$ and explicitly on a beth-fixed point $\kappa$. We prove the following. [^8] See Corollaries 4.9, 6.8, and 7.7.
@@
    For some inaccessible $\kappa$, Player II has a winning strategy in $\Game^{\mathrm{p}}_{\kappa}(o)$ (resp. $\Game^{\mathrm{sp}}_{\kappa}(o)$, $\Game^{\mathrm{ssp}}_{\kappa}(o)$). ∎
This theorem essentially characterizes when a model to a $\boldsymbol{\Sigma}_{1}$ formula can be added by forcing of the given kind. Let us mention some examples. There is a formula whose models (in all generic extensions) naturally code surjections $\omega\twoheadrightarrow\omega_{1}^{V}$. This is an example of a consistent formula which is not ssp-consistent. On the other hand, we will see a formula whose models naturally code a cofinal map $\omega\longrightarrow\omega_{2}^{V}$. [^9] See Example 2.8. This is an example of an ssp-consistent formula which is not proper-consistent. [^10] The fact that the formula is ssp-consistent follows from the fact that the Namba forcing is always ssp. We prove it explicitly in Example 8.8, but by assuming more than $\mathsf{ZFC}$. On the other hand, it is a matter of routine to verify that the cofinality of $\omega_{2}^{V}$ cannot be changed to $\omega$ by a proper poset, which explains why the formula is not proper consistent. An example of a proper-consistent formula drawn from the literature is given in Section 5.111111The formula is given implicitly, since we work with the notion of a problem. This notion and its exact relationship to infinitary propositional formulas is described in Section 2.
In addition to simply answering the question of whether a model for a given formula can be forced by a poset of a certain kind, our method provides such a poset whenever the answer is positive. The poset thus obtained is made up of conditions consisting of working parts and side-conditions. This is exactly the form of proper posets that dominated the literature on applications of $\mathsf{PFA}$ [^12] See, for example, [ACS07] or [Moo06]. Thus, our result can also be seen as an abstract unification theorem for those seemingly ad hoc constructions.
Finally, in a recent breakthrough of D. Asperó and R. Schindler,131313[AS21] a new sophisticated construction of ssp posets was presented. We observe that the concrete aim towards which their poset was built can be expressed by an infinitary propositional formula. One is then confronted by the question of whether that formula is ssp-consistent. Once this is verified, our method provides an ssp poset forcing a model for that formula.141414 Let us note here that the Asperó-Schindler poset is somewhat analogous to the posets made up of pairs of working parts and side conditions, but it did not really fit into this paradigm, while our poset is exactly of this form. Thus, the final question that we want to answer is how to verify the ssp-consistency for the formulas of “Asperó-Schindler type”. [^15] For example, in addition to the construction of [AS21] itself, its variation plays an important role in the argument of [Lie23].
We resolve this final question by providing an abstract condition on infinitary propositional formulas that we call the AS-goodness and show that it implies ssp-consistency. This condition is stated in terms of generic elementary embeddings and we will verify that the concrete formula relevant for the Asperó-Schindler result satisfies it. This verification is also an explanation of why we expect this condition to hold in other applications of their method [^16] Such as, for example, the one that appears in [Lie23]..
Let us give a brief outline of the paper.
Sec. 2
We define a simply combinatorial notion of a problem and explain why it is same whether we work with infinitary propositional formulas or with problems. We will then, for the rest of the paper, mostly work with problems, which is essentially just a matter of simplifying some proofs.
Sec. 3
We recall the characterization of genericity and semigenericity which will play the main role in designing our game-theoretic characterizations. Additionally, we introduce the notion of a virtual model and some basic notation around it.
Sec. 4
This is where we present the game-theoretic characterization of proper consistency.
Sec. 5
Here we give an example a formula whose proper consistency has been (implicitly) considered in the literature. We illustrate how to apply our method in this case.
Sec. 6
This is where we present the variation of our method that concerns the semi-proper consistency.
Sec. 7
Here we give the variation of the method for the ssp consistency.
Sec. 8
In this section, we prove the ssp-consistency criterion based on generic embeddings.
Sec. 9
Finally, we illustrate that criterion on the example of a formula coding the main task of [AS21].
@@ 3 Some Simple Remarks
The side conditions in our posets will be virtual models. They were first defined by Veličković when studying alternative ways to iterate semiproper posets, see for example [KV23]. In the cited paper, one can also find a detailed elaboration on how to work with virtual models. We review here only the basics. A prototype of a virtual model is an elementary submodel of some $H_{\theta}$, with $\theta$ regular, but there is more to the notion.
@@
If $M\in\mathcal{C}$ and if $\mathbb{Q}$ is a poset that belongs to $M\cap V_{\lambda_{M}}$, we can make sense of the phrase “$\mathbb{Q}$ is (semi)proper for $M$”. One way to do so is to just run the usual definition, but with respect to this more general object $M$212121 The point is that $M$ is not necessarily an elementary submodel of some $H_{\theta}$. It turns out that this notion is equivalent to saying that $\mathbb{Q}$ is (semi)proper for $M\cap H_{\theta}$ for any and all regular $\theta\in(2^{2^{\kappa}},\lambda_{M})\cap M$. However, $M\cap H_{\theta}\prec H_{\theta}$, so we do not diverge much from the usual notion of (semi)properness.
@@
When passing from a virtual model $M$ to its projection $M{\downarrow}\lambda$, we lose information. In a sense, $M$ then certifies that $M{\downarrow}\lambda$ is “consistent” with more information.222222 Our intuition is that it is more complicated to embed a model into a transitive structure of a larger rank. This is not true for all models $N\in\mathcal{C}_{\lambda}$ and even if it is true for some such $N$, a virtual model $M$ such that $M{\downarrow}\lambda=N$ is not necessarily unique. We will use the following notion in order to talk about this possibility of increasing the information content of a virtual model.
@@ II $q_{-1}$ $q_{0}$
Suppose that there is a proper class of inaccessible cardinals. Then the following are equivalent.
@@ 5 An Example for the Proper Consistency
The goal is to show that there exists a proper poset $\mathbb{Q}$ such that in $V^{\mathbb{Q}}$, there exists a sequence $(X_{n}:n<\omega)$ such that $X=\bigsqcup_{n<\omega}X_{n}$ and such that for all $n<\omega$, $\Gamma{\restriction}[X_{n}]^{2}=\mathrm{const}$242424 This was previously known by the results of [ARS85]. This goal can be naturally reformulated as a problem (in our formal sense).
@@
    $\phi(x)\in\mathsf{tp}^{\mathcal{H}}(a/M)$.252525 We use the notation “$\mathsf{tp}^{\mathcal{H}}(a/M)$” to denote the type of the object $a$ in the structure $\mathcal{H}$ with respect to parameters in $M$.
@@
      the first $k$ elements of $\mathsf{dom}(w_{q})$ are exactly $a_{0},\dots,a_{k-1}$,262626 We mean the first $k$ in the order $<_{\mathcal{H}}$.
@@
6 Semiproper Consistency [^27] A version of this section was done jointly with Ben De Bondt. He pursued some of these ideas in a different direction in his thesis [DB24].
@@ II $q_{-1}$ $q_{0}$
Suppose that there is a proper class of inaccessible cardinals. Then the following are equivalent.
@@ 7 Stationary Set Preserving Consistency
In this section, we want to characterize the ssp consistency of $(\mathbb{H},\mathcal{D})$ in a way analogous to what we did for proper and semiproper consistency. We follow closely the outline of Sections 4 and 6. In this case, the preconditions and their ordering are exactly the same as in the semiproper case. The game $\Game^{\mathrm{ssp}}_{\kappa}(p)$ that we define below differs from the game $\Game^{\mathrm{sp}}_{\kappa}(p)$ in the questions of the second type that Player I can ask. While a semiproper poset is semiproper for club many models, this is not true for an ssp poset. What we have is that an ssp poset is semiproper for projectively stationary many models (cf. [FJZ03, Lemma 4.8]). The change in the game is such that it reflects this fact.
@@ II $q_{-1}$ $q_{0}$
Suppose that there is a proper class of inaccessible cardinals. Then the following are equivalent.
@@ II $p_{-1}$ $p_{0}$ $p_{1}$cdots
The infinite plays with no rules broken are won by Player II. The set $\mathbb{C}^{\mathrm{BV}}_{\eta}$ is defined to consist of all $p\in\mathbb{P}^{\mathrm{sp}}\cap V_{\eta}$ such that Player II has a winning strategy in $\Game^{\mathrm{BV}}_{\eta}(p)$.282828 This definition is essentially due to De Bondt and Veličković.
@@ 9 An Example for the AS Goodness
In this section, we show how our methods can be used to give another proof of the main lemma of [AS21]. Before we precisely state what we are doing, we need a following auxiliary notion.
@@
This is shown as [AS21, Lemma 3.6]. ∎
@@
  • $M{\downarrow}\lambda$, 9
  • $\Game^{\mathrm{BV}}_{\eta}(p)$, 41
@@
  • $\Game^{\mathrm{p}}_{\kappa}(p)$, 11
  • $\Game^{\mathrm{sp}}_{\kappa}(p)$, 27
  • $\Game^{\mathrm{ssp}}_{\kappa}(p)$, 34
@@
  • $\beth_{\mathrm{fix}}$, 9
  • $\delta(M)$, 8
  • $\delta_{M}$, 8
  • $\lambda_{M}$, 9
  • $\leq^{\mathrm{p}}$, 11
  • $\leq^{\mathrm{sp}}$, 27
  • $\mathbb{C}^{\mathrm{BV}}_{\eta}$, 41
  • $\mathbb{C}^{\mathrm{p}}_{\kappa}$, 12
  • $\mathbb{C}^{\mathrm{sp}}_{\kappa}$, 28
  • $\mathbb{C}^{\mathrm{ssp}}_{\kappa}$, 35
@@
  • $\mathbb{P}^{\mathrm{sp}}$, 27
  • $\mathcal{C}$, 9
  • $\mathcal{C}_{\lambda}$, 9
@@
  • $\mathsf{Hull}(M,X)$, 9
  • $\phi_{\mathrm{AS}}$, 51
  • $\theta$-well behaved, 48
  • $\vartheta(\mathcal{M})$, 49
  • $o$, 12
  • $pq$, 6
  • AS-good, 40
@@
  • problem, 6
  • proper-consistent, 2, 6
  • semiproper-consistent, 2, 6
  • ssp, 2
  • ssp-consistent, 2, 6
  • virtual model, 9
  • vm-chain, 10